OpenStudy (anonymous):

Dy/dx=x(lny). y(0)=1

OpenStudy (anonymous):

Is it \(\ln(y(x))\) or \(\ln(x)\)?

OpenStudy (anonymous):

No it's just like I wrote it I know I would do this Dy/lny=xdx but then what

OpenStudy (anonymous):

Integrate both sides. Here you would receive: \[ \int \frac{1}{\ln(y)}dy=\int x\,dx\implies\\ \text{Li}(y(x))=\frac{x^2}{2}+C \implies\\ y(x)=\text{Li}^{-1}\left(\frac{x^2}{2}+C\right) \]

OpenStudy (anonymous):

Wait, change up the \(\text{Li}(x)\) to \(\text{li}(x)\)

OpenStudy (anonymous):

(Sorry, I mixed up my offset log int)

OpenStudy (anonymous):

Huh? Dy/lny

OpenStudy (anonymous):

Yep. Integrating that gives you a function called the logarithmic integral, here: en.wikipedia.org/wiki/Logarithmic_integral_function

OpenStudy (anonymous):

Nope! It's not as hard as that. Consider the behavior of the function in the neighborhood of the point \((0,1)\) (the initial value). \(\ln(1)=0\), so \(y'=0\). This means that in this neighborhood the function is approximately stationary. But, if there is no change in \(y\), then \(y=1\) and the function is stationary globally. By the uniqueness theorem, we must conclude that \(y(x)=1\) is the only solution. No special functions needed!

OpenStudy (anonymous):

Check the answer! if \(y(x)=1\), \(y'(0)=0\) \(x\ln(y(x))=x\ln(1)=0\)

OpenStudy (anonymous):

@kmalone99